OpenStudy (anonymous):

7x - 3/8 = 6x - 5/8 please help

OpenStudy (lilg132):

3/8 = 0.375 5/8 = 0.625 7x - 0.375 = 6x - 0.625 7x - 6x = 0.375 - 0.625 x = -0.25 or x = -1/4

OpenStudy (lilg132):

remember to click good answer

OpenStudy (anonymous):

lol thank u